4. Amal Automotive Company experienced the following accounting events during 2019. Received S60,000 cash from the issue of common stock. Performed services for $35,000 cash. Purchase land for $16,000 cash. Borrowed $1,000 cash from the Bank. Paid $24,000 cash for salary expenses. Sold land for $19,000 cash. Paid $20,000 cash on the loan from state bank. Paid $12,800 cash for salary expense. Paid a cash dividend of $15,000 to the stockholders Required: a. Pass general journal entries and post the entries, to T-accounts and determine the ending balance in each account. b. Prepare a statement of cash flows. Assume Amal Automotive had a beginning cash balance of $10,000.
1
Expert's answer
2020-03-25T10:34:05-0400

a) beginning cash balance of $10,000

debit Turnovers: 60 000+35 000+19 000+1000 =115 000

credit Turnovers:16 000+ 24 000+20 000+12 800+ 15 000=87 800

Balance at the end: 10 000+115 000-87 800= 37 200

general journal entries:

Debit cash Credit issue of common stock - 60 000

Debit cash Credit Performed services - 35 000

Debit cash Credit Bank - 1 000

Debit cash Credit land buyer- 19 000

Debit land seller Credit cash - 16 000

Debit salary expenses Credit cash - 24 000

Debit  loan from state bank Credit cash - 20 000

Debit salary expenses Credit cash - 12 800

Debit stockholders Credit cash - 15 000

b) Operating Activities: 35 000-12 800-24 000=-1800

Investing Activities: 19 000-16 000= 3 000

Financing Activities: 60 000+1000-20 000-15 000=26 000

Total: 10 000+26 000+3 000-1 800=37 200